# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v1

@@ Appendix F Non-existence of continuous dynamical invariant
Without loss of generality, assume $y\geq 0$. Now fix any point $\theta\in T^{-1}(\lambda-2/\eta,4/\eta-2(y+\lambda))$. Under the conjugacy (7), we have that $T(\theta)=(\nu-2,4-2(\mu+\nu))$. Then according to Proposition 25 for the regularized case and Proposition 20, in any neighborhood $O$ of $T(\theta)$, there exists $\xi^{\prime},\xi^{\prime\prime}$ and $N^{\prime},N^{\prime\prime}$ such that $F^{N^{\prime}}(\xi^{\prime})=T(\theta^{\prime})$ and $F^{N^{\prime\prime}}(\xi^{\prime\prime})=T(\theta^{\prime\prime})$. Using the same argument as in Appendix E.2.2 and in Appendix E.1.2, we have that there exists $\bar{\theta}^{\prime},\bar{\theta}^{\prime\prime}$ such that $T(\mathrm{GD}_{\eta}^{N^{\prime}}(\bar{\theta}^{\prime}))=T(\theta^{\prime})$ and $T(\mathrm{GD}_{\eta}^{N^{\prime\prime}}(\bar{\theta}^{\prime\prime}))=T(\theta^{\prime\prime})$. Next, we show that $\bar{\theta}^{\prime}$ and $\bar{\theta}^{\prime\prime}$ can be chosen such that $\mathrm{GD}^{N^{\prime}}_{\eta}(\bar{\theta}^{\prime})=\theta^{\prime}$ and $\mathrm{GD}^{N^{\prime\prime}}_{\eta}(\bar{\theta}^{\prime\prime})=\theta^{\prime\prime}$. To see this, notice that, for any $\in\mathbb{R}^{2}, (u,v)=(s,t)(u,v)=(s,t)$ if and only if $T(u,v)=T(s,t)$ and the two pairs, $u+v$ and $s+t$, and, uvu-v and sys-y, have the same sign. Consider change of coordinate $p=(u+v)/\sqrt{2}$ and $q=(u-v)/\sqrt{2}$. Let $p_{k}=(u_{k},v_{k})$ denote the orbit under $\mathrm{GD}_{\eta}$. By direct computation, we have that
@@
Therefore, the sign of $p_{N^{\prime}}$ is fully determined by whether number of $z_{j}$ such that $1-\nu>z_{j}$ is even or odd. We denote this number by $n_{p}$. Similarly, we have
@@
and the sign of $q_{N^{\prime}}$ is fully determined by whether number of $z_{j}$ such that $z_{j}>\nu-1$ is even or odd. We denote this number by $n_{q}$. Notice that we can take $\xi^{\prime}$ as follows
$\xi^{\prime}=G_{0}^{m_{p}}\circ G_{2}^{m_{q}}(T(\theta^{\prime})),$
here $m_{q}\in\left\{0,1\right\}$, and $m_{p}$ is any large enough integer. Note since the image of $G_{0}$ lies out side $\left\{z>1-\nu\right\}$, $m_{p}$ does not have an effect on $n_{q}$. Also, since the image of $G_{2}$ is contained in $\left\{z>1-\nu\right\}$, one can always select $m_{q}$ from $\left\{0,1\right\}$ to make $n_{q}$ even or odd. Meanwhile, since the image of $G_{0}$ is contained in $\left\{z<\nu-1\right\}$, one can always select a $m_{p}$ to make $n_{p}$ even or odd. Therefore, the sign of $p_{N^{\prime}}$ and $q_{N^{\prime}}$ can be arbitrary. This implies that, one can always select $\bar{\theta}^{\prime}$ such that $\mathrm{GD}^{N^{\prime}}_{\eta}(\bar{\theta}^{\prime})=\theta^{\prime}$. A similar statement holds for $\bar{\theta}^{\prime\prime}$.
@@ Appendix G General matrix factorization
The gradient descent update map $\mathrm{GD}_{\eta}$ is non-invertible in general. Nevertheless, we show that the parameter space can be partitioned into small pieces, so that confined on each piece, $\mathrm{GD}_{\eta}$ has a simple behavior.
@@ Appendix H Experiment details
For Figure 2, we consider $L(x,y)=(xy-1)^{2}$ with $(x,y)\in\mathbb{R}^{2}$. For the left panel, we evenly sampled $800^{2}$ initial points in the range $4.5,4.5]^{2}$ and ran gradient descent with step size $0.2$ for $6$ iterations. To visualize the basin for unstable minimizers, note, as shown in the proof of Proposition 15, converging to unstable minimizers can only occur within finitely many steps. We therefore recorded the final loss value and used the “contour” function from the Matplotlib package (version $3.5.2$) to collect points in the level set of 0 for the loss. Those points correspond to convergence to a global minimizer within $6$ or less steps. We then filtered out and visualized points that converge to an unstable global minimizer, i.e., a minimizer with squared norm larger than $2/\eta$ (see Corollary 16).

### v2

@@ Appendix F Non-existence of continuous dynamical invariant
Without loss of generality, assume $y\geq 0$. Now fix any point $\theta\in T^{-1}(\lambda-2/\eta,4/\eta-2(y+\lambda))$. Under the conjugacy (7), we have that $T(\theta)=(\nu-2,4-2(\mu+\nu))$. Then according to Proposition 25 for the regularized case and Proposition 20 for the unregularized case, in any neighborhood $O$ of $T(\theta)$, there exists $\xi^{\prime},\xi^{\prime\prime}$ and $N^{\prime},N^{\prime\prime}$ such that $F^{N^{\prime}}(\xi^{\prime})=T(\theta^{\prime})$ and $F^{N^{\prime\prime}}(\xi^{\prime\prime})=T(\theta^{\prime\prime})$. Using similar arguments as those in Appendix E.2.2 and in Appendix E.1.2, we have that there exists $\bar{\theta}^{\prime},\bar{\theta}^{\prime\prime}$ such that $T(\mathrm{GD}_{\eta}^{N^{\prime}}(\bar{\theta}^{\prime}))=T(\theta^{\prime})$ and $T(\mathrm{GD}_{\eta}^{N^{\prime\prime}}(\bar{\theta}^{\prime\prime}))=T(\theta^{\prime\prime})$. Next, we show that $\bar{\theta}^{\prime}$ and $\bar{\theta}^{\prime\prime}$ can be chosen such that $\mathrm{GD}^{N^{\prime}}_{\eta}(\bar{\theta}^{\prime})=\theta^{\prime}$ and $\mathrm{GD}^{N^{\prime\prime}}_{\eta}(\bar{\theta}^{\prime\prime})=\theta^{\prime\prime}$. To see this, notice that, for any $\in\mathbb{R}^{2}, (u,v)=(s,t)(u,v)=(s,t)$ if and only if $T(u,v)=T(s,t)$ and, the two pairs, $u+v$ and $s+t$, and, uvu-v and sts-t, have the same sign. Consider the change of coordinates $p=(u+v)/\sqrt{2}$ and $q=(u-v)/\sqrt{2}$. Let $(u_{k},v_{k})_{k\geq 0}$ denote an orbit under $\mathrm{GD}_{\eta}$. By direct computation, we have that
@@
Therefore, the sign of $p_{N^{\prime}}$ is fully determined by whether
$n_{p}=\#\left\{j\in\left\{0,\cdots,N^{\prime}-1\right\}\colon 1-\nu<z_{j}\right\}$
is even or odd. Similarly, we have
@@
and the sign of $q_{N^{\prime}}$ is fully determined by whether
$n_{q}=\#\left\{j\in\left\{0,\cdots,N^{\prime}-1\right\}\colon 1-\nu>z_{j}\right\}$
is even or odd. Notice that we can take $\xi^{\prime}$ as follows
$\xi^{\prime}=G_{0}^{m_{q}}\circ G_{2}^{m_{p}}(T(\theta^{\prime})),$
where $G_{0}$ and $G_{2}$ are defined as in Proposition 13, and $m_{p}\in\left\{0,1\right\}$, and $m_{q}$ is a sufficiently large number. Since the image of $G_{0}$ lies out side $\left\{z>1-\nu\right\}$, increasing $m_{q}$ does not affect $n_{p}$. Also, since the image of $G_{2}$ is contained in $\left\{z>1-\nu\right\}$, one can always select $m_{p}$ from $\left\{0,1\right\}$ to make $n_{p}$ even or odd as needed. Now fix $m_{p}$. Since the image of $G_{0}$ is contained in $\left\{z<\nu-1\right\}$, one can always select a sufficiently large $m_{q}$ to make $n_{q}$ even or odd as needed. Consequently, by appropriate choices of $m_{p}$ and $m_{q}$, the signs of $p_{N^{\prime}}$ and $q_{N^{\prime}}$ can be made arbitrary. This implies that, one can always select $\bar{\theta}^{\prime}$ such that $\mathrm{GD}^{N^{\prime}}_{\eta}(\bar{\theta}^{\prime})=\theta^{\prime}$. A similar statement holds for $\bar{\theta}^{\prime\prime}$.
@@ Appendix G General matrix factorization
The gradient descent update map $\mathrm{GD}_{\eta}$ is non-invertible in general. Nevertheless, we show that the parameter space can be partitioned into small pieces, so that when restricted to each piece, $\mathrm{GD}_{\eta}$ has a simple behavior.
@@ Appendix H Experiment details
For Figure 2, we consider $L(x,y)=(xy-1)^{2}$ with $(x,y)\in\mathbb{R}^{2}$. For the left panel, we evenly sampled $800^{2}$ initial points in the range $4.5,4.5]^{2}$ and ran gradient descent with step size $0.2$ for $6$ iterations. To visualize the basin for unstable minimizers, note, as shown in Corollary 16, converging to unstable minimizers can only occur within finitely many steps. We therefore recorded the final loss value and used the “contour” function from the Matplotlib package (version $3.5.2$) to collect points in the level set of 0 for the loss. Those points correspond to convergence to a global minimizer within $6$ or less steps. We then filtered out and visualized points that converge to an unstable global minimizer, i.e., a minimizer with squared norm larger than $2/\eta$ (see Corollary 16).